### v1

# LabUtopia: High-Fidelity Simulation and Hierarchical Benchmark for Scientific Embodied Agents

Abstract
Scientific embodied agents play a crucial role in modern laboratories by automating complex experimental workflows. Compared to typical household environments, laboratory settings impose significantly higher demands on perception of physical-chemical transformations and long-horizon planning, making them an ideal testbed for advancing embodied intelligence. However, its development has been long hampered by the lack of suitable simulator and benchmarks. In this paper, we address this gap by introducing LabUtopia, a comprehensive simulation and benchmarking suite designed to facilitate the development of generalizable, reasoning-capable embodied agents in laboratory settings. Specifically, it integrates i) LabSim, a high-fidelity simulator supporting multi-physics and chemically meaningful interactions; ii) LabScene, a scalable procedural generator for diverse scientific scenes; and iii) LabBench, a hierarchical benchmark spanning five levels of complexity from atomic actions to long-horizon mobile manipulation. LabUtopia supports 30 distinct tasks and includes more than 200 scene and instrument assets, enabling large-scale training and principled evaluation in high-complexity environments. We demonstrate that LabUtopia offers a powerful platform for advancing the integration of perception, planning, and control in scientific-purpose agents and provides a rigorous testbed for exploring the practical capabilities and generalization limits of embodied intelligence in future research.
Figure 1: The LabUtopia simulation environment and benchmark for developing scientific embodied agents in automated laboratories. LabUtopia supports chemical reaction modeling and provides diverse laboratory assets, forming a high-fidelity testbed for tasks of varying difficulty—from atomic actions to long-horizon action sequences involving both manipulation and navigation.
1 Introduction
Scientific breakthroughs play a foundational role in advancing human knowledge shah2023foundation , driving technological innovation, and improving societal well-being anadon2016making . However, the traditional paradigm of natural science research remains slow and labor-intensive shen2021fourth , where countless experiments must be performed by skilled researchers to reach meaningful insights takahashi2007induction ; nicolaou1994total ; onuma1999identification ; bednorz1986possible . These limitations constrain the overall pace and scalability of scientific discovery. Therefore, automated laboratories dai2024autonomous ; jiang2022artificial have emerged as a promising alternative, aiming at developing intelligent agents that autonomously design and execute complex experiments through adaptive workflows. By reducing human workload, such agents enable scalable, reproducible, and around-the-clock experimentation, significantly increasing research throughput hickman2023self . Nevertheless, building effective scientific agents remains challenging, particularly due to the high cost of data collection and the difficulty of generalizing across diverse hardware platforms musslick2025automating .
A promising approach to address the above issues is the sim-to-real framework zhao2020sim , where agents are first trained within realistic simulations before being deployed in real-world laboratory settings. This paradigm enables cost-effective and safe training while maintaining the potential for generalization to physical environments hsu2023sim . However, current established simulators predominantly focus on household environments kolve2017ai2 ; puig2018virtualhome ; yang2024holodeck and fail to adequately address the specific challenges of scientific experimentation. As discussed in Table 1, they exhibit three fundamental limitations: (1) the inability to model chemical dynamics, such as product formation or color changes, which are essential for accurate perception and reasoning in lab tasks; (2) limited diversity and semantic richness in their asset libraries, which restricts the faithful representation of heterogeneous laboratory environments in real world; and (3) the lack of comprehensive evaluation protocols, particularly those that span from fine-grained atomic actions to complex, long-horizon experimental procedures.
To address these challenges, we present LabUtopia, a comprehensive simulation and benchmarking suite tailored for scientific laboratory contexts. LabUtopia integrates a diverse set of functional assets, a hierarchical task taxonomy, and a high-fidelity simulation engine capable of modeling rigid, deformable, and fluid objects, as well as simulating both physical and chemical processes. Our goal is to provide a scalable, versatile platform for training and evaluating agents’ perception, planning, and control skills under high task complexity and diverse environments, advancing the role of embodied intelligence in accelerating scientific discovery. The key designs and innovations include:
(1) LabSim is a high-fidelity simulation environment built on Isaac Sim, enhanced with a chemical engine that models reaction-driven transformations (e.g., color change, product generation) by combining a curated substance database with the reasoning model. Extending beyond conventional physical dynamics, LabSim supports a wide range of chemical reactions, enabling precise and visually grounded simulation of laboratory phenomena.
(2) LabScene is a procedural generation pipeline that synthesizes diverse, physically plausible 3D laboratory scenes aligned with real-world configurations. Built upon a curated set of expert-verified assets, LabScene employs a hybrid layout strategy that combines grid stochastic sampling with constraint-aware search, enabling scalable environment creation for scientific embodied tasks.
(3) LabBench is a hierarchical benchmark, featuring a five-level task structure that spans from atomic object manipulations to long-horizon missions requiring integrated navigation and manipulation. Together, these components establish a rigorous and scalable testbed for developing and evaluating scientific-purpose embodied agents under rich physical constraints and procedural complexity.
To provide an in-depth analysis of embodied agents in scientific laboratory settings, we complement prior research with an evaluation that targets the agent’s ability to conduct realistic experimental procedures that involve accurate material recognition, multi-step task planning, and precise instrument control. Powered by our LabScene generation pipeline, we construct over 100 diverse and physically plausible lab environments, and evaluate agent performance across 30 tasks of varying complexity in the LabBench benchmark. Through extensive experiments, we show that current state-of-the-art manipulation policy models still struggle with the variability of instrument configurations and accumulated errors in long-horizon task execution, highlighting the need for more specialized solutions in research-oriented embodied AI.
We summarize our main contributions as follows:
  • We present LabUtopia, a simulation and benchmarking suite tailored to the unique challenges of laboratory settings. LabUtopia supports complex physical interactions and various chemical reactions across 30 distinct tasks, enabling realistic evaluation of embodied agents in high-fidelity scientific scenarios.
  • We provide a high-quality asset set comprising over 100 laboratory scenes and 100 scientific instruments, that have been standardized and filtered by domain experts. Building on this, we design an automated scene generation pipeline to produce diverse, scalable lab environments, supporting both real-world alignment and large-scale training and evaluation.
  • We introduce a hierarchical benchmark that spans multiple levels of task complexity, from low-level atomic operations to high-level long-horizon reasoning tasks. This structure enables principled assessment of embodied agents’ capabilities and reveals performance bottlenecks across varying levels.
2 Related Work
Automated Laboratories. Current automated laboratories enhance the efficiency and scalability of experiments in chemistry and materials science while reducing costs by integrating machine learning, robotics and modular platforms. Self-driving laboratories (SDLs) abolhasani2023rise automate repetitive tasks, yet often constrain autonomous experimental design. Systems such as Synbot ha2023ai , Chemputer steiner2019organic , MARS-Chem dai2024autonomous , Artificial Chemist abdel2021self , Reactivity Explorer dragone2017autonomous , and AI-EDISON jiang2022artificial have demonstrated remarkable performance in organic synthesis, exploratory chemistry, and nanomaterial optimization, enabling standardized and reproducible experiments and accelerating molecular discovery. However, these systems are limited by predefined protocols, hardware dependencies, insufficient task comprehension, and poor real-time adaptability. As a result, their flexibility and intelligence are constrained, making them less suitable for various experimental tasks. Moreover, most of these systems primarily focus on advancing scientific discovery, with limited attention to the long-term development of intelligent embodied systems, thereby overlooking the potential value of embodiment in scientific research. Therefore, we propose a LabUtopia that offers low-cost, high-efficiency workflows and large-scale data acquisition, enhancing flexibility and data-driven capabilities. This environment aims to support the development of embodied intelligence that is adaptable to a wide range of chemical research scenarios.
Simulators for Embodied AI. The rapid development of simulators is progressively transitioning from general-purpose functionality to high-fidelity realism. Certain simulators emphasize versatile capabilities, primarily for modeling interactions and dynamic changes in the physical world, facilitating algorithm validation and training. PyBullet coumans2016pybullet , Gazebo koenig2004design , and RLBench james2020rlbench support real-time physical simulations, including rigid body dynamics and collision detection, while offering diverse sensor emulation and deep learning integration. Conversely, other simulators prioritize high-fidelity scene reconstruction to meet the demands of complex real-world task environments. ARNOLD gong2023arnold , VLMbench zheng2022vlmbench , Habitat savva2019habitat , OmniGibson li2023behavior ; srivastava2022behavior , ManiSkill3 tao2024maniskill3 , and ClevrSkills haresh2024clevrskills focus on language-guided task learning in realistic 3D environments, aiming to advance robotic manipulation and human-robot interaction research. These platforms offer vision-language manipulation benchmarks, open-source frameworks, and human-centric evaluations, supporting rich simulations of daily activities, GPU-accelerated parallel robot simulation, and photorealistic rendering, while investigating compositional reasoning and generalization capabilities. However, existing simulation platforms generally lack specialized modeling for laboratory settings and operations. To address this, we propose LabUtopia based on Isaac Sim nvidia2023isaacsim tailored for chemical laboratories, enabling embodied agents to perform operational learning, path navigation, and task planning in chemical experimental environments. Integrated with visualized simulations of chemical reaction processes, LabUtopia aims to provide critical support for the advancement of embodied intelligence in experimental sciences.
Simulator/Benchmark Simulation Lab Environment Task Diversity
Fluid Physics Chemistry Scene Object Multi-action Composed Generalization Long-Horizon Navigation
Alfred shridhar2020alfred
Behavior li2023behavior ; srivastava2022behavior
RLBench james2020rlbench
Ravens zeng2021transporter
Vlmbench zheng2022vlmbench
Arnold gong2023arnold
VIMA-Bench jiang2022vima
Maniskill3 tao2024maniskill3
ClevrSkills haresh2024clevrskills
LabUtopia (Ours)
Table 1: Comparison with existing embodied AI simulators/benchmarks. Fluid / Physics / Chemistry: Support for simulating fluids, realistic physical interactions, and chemical processes, respectively. Scene / Object: Whether the benchmark supports realistic lab environments with scene-level and object-level assets. Multi-action: Multiple different axiom actions. Composed: Tasks involve the simple composition of atomic actions. Generalization: Generalization task across unseen scenes or object variations. Long-Horizon: Tasks demand high-level planning and long sequences of atomic and composed actions. Navigation: Tasks integrate spatial navigation with manipulations.
3 Laboratory Simulation Suite
We introduce LabUtopia, a high-fidelity simulation platform tailored to the challenges of embodied manipulation in laboratory settings. It is specifically designed for simulating, training, and evaluating agents in lab-centric tasks. LabUtopia consists of three key components: LabSim provides high-fidelity physical simulation with extensions for modeling chemically relevant dynamics, such as fluid mixing and reactive state transitions. LabScene includes a diverse asset library of scientific instruments and procedurally generates 3D environments, enabling rich spatial and task variations. Finally, a built-in trajectory collection module supports automated generation of expert demonstrations, facilitating scalable data collection for diverse lab tasks.
3.1 LabSim: High-Fidelity Simulation Environment
LabSim is a high-fidelity simulation engine designed to model the rich physical and chemical phenomena of laboratory environments. It not only supports physically accurate modeling of diverse material properties, but also introduces a reasoning-driven pipeline to simulate chemical reactions, enabling aligned training and evaluation with real-world scientific workflows.
Physical Realism. LabSim supports physically accurate interactions among rigid, deformable, and fluid entities PhysX5 . Each asset in the environment is annotated with empirically grounded physical properties such as mass, friction coefficients, and restitution. For rigid-body simulation, we enable precise contact and collision modeling. For soft materials, we incorporate deformable body physics to capture compressible and elastic behavior. Notably, for fluid simulation, we employ a GPU-accelerated Position-Based Dynamics (PBD) framework macklin2013position , supporting rich fluid-agent interactions required for scientific manipulation.
Chemical Process Modeling. To simulate chemical processes within laboratory tasks, we introduce a chemical engine that integrates a curated knowledge base with a reasoning model. We begin by constructing a structured database of 200 common chemical substances, sourced from the authoritative PubChem repository kim2016pubchem . Each encodes key its attributes, such as color, molar mass, and pH value, allowing it to be represented as a substance asset within the simulation. Given a set of reactants, we leverage a large language model (GPT-4o mini hurst2024gpt ) to reason about potential chemical processes and infer corresponding transformations, including color changes, product formation, etc. These inferred changes are then rendered in the simulation by dynamically updating the physical state and visual properties of the involved substances. This engine equips LabUtopia with the capability to model complex chemical interactions with both interpretability and flexibility.
3.2 LabScene: Procedural Laboratory Environment Generation at Scale
Current 3D scene datasets primarily focus on domestic, office, or industrial environments li2021igibson ; tao2024maniskill3 ; baruch2021arkitscenes ; fu20213d ; hua2016scenenn , offering limited support for simulating laboratory settings. However, training and evaluating embodied agents in lab-centric tasks requires high-quality, interactive environments populated with scientifically relevant instruments and layouts ir.2025.11 . To address this gap, we introduce LabScene, a scalable dataset of laboratory object and scene assets with the procedural generation mechanism.
Figure 2: An overview of our laboratory simulation suite. LabScene automatically synthesizes scalable laboratory scenes using a diverse asset library and a procedural generation pipeline, while LabSim supports the simulation of high-fidelity physical and chemical interactions.
Scene Assets. Due to the scarcity of open-source lab assets in the community, we collected over one thousand candidate scenes from designer websites. These raw assets underwent a multi-stage pre-processing pipeline involving content filtering, format normalization, and structural standardization. To ensure realism, we consulted domain experts in chemistry and physics to assess scene fidelity and provide refinement suggestions. Based on their feedback, we selected a high-quality subset of expert-verified scenes to serve as the foundational environments for LabUtopia.
Object Assets. While some collected scenes include basic instruments, many lack detailed internal structures required for executing lab tasks. To address this, we assembled a comprehensive set of object assets spanning essential equipment (e.g., drying ovens) and diverse glassware types (e.g., beakers). To ensure compatibility with robotic manipulation, we refine and modularize these assets into standard, interactable objects. And some were augmented with articulated joints and hinge mechanisms to support physically plausible interactions within the simulation. In total, the final dataset includes about 60 categories of laboratory equipment assets and around 80 types of transparent glassware and plasticware assets, covering a diverse range of materials, sizes, and functional forms.
Environment Generation Pipeline. To incorporate our collected instruments into the laboratory scenes, we develop an environment generation pipeline that balances layout diversity with physical plausibility. Specifically, all objects are placed sequentially based on importance and size rankings. For each object, candidate positions and orientations are sampled from a discretized grid that satisfies various constraints, including boundary, collision, and instrument-specific constraints deitke2022️ . The layout score is computed considering factors such as edge proximity, inter-object distance, and orientation alignment. The configuration with the highest score is selected. If the random sampling fails to produce a valid layout within a time limit, the system falls back to a depth-first search strategy sun2024forest2seq , systematically exploring placements while enforcing physical and spatial constraints. This hybrid approach ensures functional, reasonable scene layouts suitable for embodied agent training.
3.3 Task-Rich Trajectory Collection
Manipulation Trajectory Auto-collection. We divide the motion planners into two levels: atomic action controllers and task-level action controllers. Atomic actions are standard laboratory operations, such as pouring and stirring, that align with experimental protocols and are executed using finite state machines. Task-level controllers organize atomic action controllers to collect data for specific tasks. For atomic actions, we control multiple target keypoints for the robotic arm. At each keypoint, an RMPflow controller cheng2020rmp plans motion toward dynamically determined positions based on the real-time state of manipulated objects. We incorporate spherical linear interpolation (Slerp) gong2023arnold for continuously manipulating articulated objects, e.g., opening the dry container task, for more robust results. Furthermore, task-level controllers organize atomic actions to streamline the entire experimental procedure, enabling our motion planner to generate demonstration data efficiently.
Navigation Trajectory Auto-collection. Robots in laboratory environments need to autonomously move between locations and interact with various experimental instruments. To automate trajectory collection, we design a method based on the A* algorithm guruji2016time and occupancy map elfes2002using . Specifically, we first generate and store occupancy maps built by Isaac Sim for the laboratory. These maps explicitly indicate the areas where obstacles are present, thus marking regions that are non-navigable. This information is then bound to the laboratory asset data. When deployed in a new laboratory scene, the robot uses its initial and target positions to plan a path with key waypoints. It then follows these waypoints, generating navigation trajectory data. This approach proves successful in experiments, offering an efficient solution for trajectory planning and data collection.
Figure 3: An overview of our hierarchical benchmark. LabBench structures scientific tasks across five levels, from atomic manipulations to long-horizon experiments, enabling rigorous evaluation of embodied agents in realistic laboratory settings.
4 LabBench: Hierarchical Benchmark for Lab Agents
Embodied manipulation in laboratory environments spans a wide spectrum of tasks, ranging from low-level interactions to long-horizon workflows. These tasks vary significantly in complexity and skill requirements, making it difficult to evaluate agent capabilities in a unified manner. To address this, we propose LabBench, a hierarchical benchmark comprising over 50 tasks designed to systematically evaluate embodied agents across multiple levels of control, planning, and reasoning.
4.1 Five-Level Task Structure
In order to comprehensively evaluate embodied agents in laboratory environments, LabBench organizes its tasks into five levels of increasing complexity, ranging from low-level atomic actions to integrated long-horizon and mobile manipulation tasks.
Level 1: Atomic Manipulation Tasks. This level focuses on fundamental low-level interactions that serve as the building blocks for more complex operations. Tasks include single-step action such as grasping, pouring, stirring, opening instrument, and placing containers. These actions can typically be executed via primitive controllers without requiring task-level planning.
Level 2: Short-Horizon Manipulation Tasks. This level involves agents performing a sequence of 2-3 atomic actions to complete a compound objective. For example, an agent might open a container and then pour a reagent, or pick up a test tube and place it in a mixer. These tasks require precise coordination of sequential actions to achieve the desired outcome.
Level 3: Generalizable Short Manipulation Tasks. This level evaluates agents’ generalization capabilities under distributional shifts. Agents are trained jointly on mix of objects with varying shapes and appearances, as well as different visual material and environmental. Tasks are tested in novel scenarios featuring unseen object configurations, unfamiliar scene arrangements, or appearance variations. The evaluation tests the agents’ capacity to transfer learned skills to effectively handle out-of-domain objects and scenes.
Level 4: Long-Horizon Manipulation Tasks. This level involves high-level planning and executing multi-step laboratory protocols that span numerous atomic and composed actions. These workflows, such as preparing a chemical solution or executing the cleaning instrument program, require high-level planning, reasoning, and robustness to compounding execution errors.
Level 5: Mobile Manipulation Tasks. The highest level integrates spatial navigation with manipulation. Agents are required to traverse large-scale laboratory environments using mobile-base control while performing manipulation tasks. Our task is to transport container between areas.
4.2 Evaluation Protocol
Embodiment. We employ a 7-DoF Franka Emika Panda FrankaRobotics2025 manipulator with a parallel gripper for manipulation tasks. The agent controls the seven joints and gripper directly. Isaac Sim’s built-in motion planner maps end-effector commands to joint-space actions for execution. For navigation tasks, we utilize the Fetch mobile manipulator FetchRobotics2025 and Ridgeback with Franka ClearpathRobotics2025 , both supporting arm manipulation and base locomotion. These robots are controlled using three degrees of freedom—x and y velocities and rotation—enabling integrated navigation and manipulation wu2020towards .
Evaluation Execution. A task instance is considered successful if the success condition—where the current state remains within the tolerance threshold of the goal state—is continuously satisfied for 2 seconds after the agent completes the final stage of the task, which is the same as previous work gong2023arnold ; he2024learning ; hu2024videopredictionpolicygeneralist . For example, in the “Open Door” task, success requires the cabinet door to remain at the specified open position for 2 seconds after the robot releases the handle, ensuring no shortcuts during motion. The success rate is used as the evaluation metric in LabSim, with strict evaluation ensuring both the robot and object maintain the successful state for the required duration, only after the motion planner has executed its final action.
5 Experiment
5.1 Experimental Setup
Models. To benchmark the performance of existing imitation learning algorithms in LabSim, we select two representative models: ACT fu2024mobile and Diffusion Policy chi2023diffusion .
• ACT is a transformer-based model designed for sequential decision making in robotic manipulation. ACT processes RGB images, robot proprioception, and tokenized instructions through a multi-layer transformer encoder. At each decision step, ACT autoregressively predicts the next low-level action, conditioned on past observations and actions.
• Diffusion Policy is a generative model that formulates robot control as a conditional diffusion process. The model takes recent observations as input and learns to generate control trajectories by progressively denoising an initial random trajectory sample, conditioned on the observation context. This approach not only enables efficient learning from expert demonstrations, but also provides robust stochasticity during action generation. In our work, we utilize the CNN-based version of Diffusion Policy as the encoder.
Table 2: Performance comparison across task levels. For Level-3, results are reported under ID/OOD settings. Values represent success rates (%).
Task Level Task Name ACT DP
Level-1 Stir 86.7 95.0
Pick 86.7 75.0
Pour 76.7 73.3
Press 93.3 96.7
Place 73.3 80.0
Shake 93.3 86.7
Open Door 77.5 67.5
Close Door 71.7 65.0
Open Drawer 93.3 95.0
Close Drawer 96.7 100.0
Level-2 Pour Liquid 67.5 50.0
Shake Beaker 77.5 67.5
Heater Beaker 86.7 25.0
Operate Drawer 73.3 10.0
Stir w/ GlassRod 55.0 10.0
Transport Beaker 78.3 50.0
Level-3 Pick 81.7 / 71.7 53.3 / 41.7
Press 98.3 / 96.7 81.7 / 31.6
Open Door 73.3 / 65.0 63.3 / 58.3
Pour Liquid 75.0 / 65.0 46.6 / 31.6
Heater Beaker 86.7 / 80.0 21.6 / 8.3
Transport Beaker 77.5 / 73.3 67.5 / 15.0
Visual Input. We use two cameras for each manipulation task, positioned to ensure both the robot and the objects are visible. Camera placement may vary slightly by task, but the general principle remains consistent. Each camera outputs 256×256 RGB images by default. Since accurate depth data for transparent objects is difficult to obtain in the real world, depth maps are not included by default, but users can render images at any resolution and optionally enable depth maps. Other Omniverse sensors (e.g., tactile sensors) are currently disabled as they are unnecessary for current tasks, but full support is available if needed.
Training Details. All experiments are conducted on a single Nvidia 4090 GPU. For each task, the collected dataset consists of 150 episodes. Evaluation is performed in real-time within a simulated environment based on the task settings, with 60 episodes used for testing. Following the original settings in ACT and Diffusion Policy, we set the horizon to 8 and 60, respectively, and the number of training epochs is uniformly set to 200.
Table 3: Performance comparison of different models on Level-4 long-horizon tasks. SP: Single-stage Primitive; A1–A7: Different sub-steps of task sequence.
Level-4 Task Model Success Rate (%)
SP A1 A2 A3 A4 A5 A6 A7
Clean Beaker ACT 14.0 99.3 51.9 43.3 42.5 12.3 10.6 1.6
DP 2.5 92.7 8.1 1.3 0.0 0.0 0.0 0.0
Drying Beakers ACT 6.3 81.3 42.0 19.7 9.3 3.5 1.2 0.0
DP 2.0 76.0 32.0 26.3 0.0 0.0 0.0 0.0
Liquid Fusion ACT 8.3 96.3 57.6 46.7 42.5 12.3 10.6 1.6
DP 6.7 93.5 12.3 8.3 0.0 0.0 0.0 0.0
5.2 Experimental Results
We evaluated the performance of the ACT and DP models across Level-1 to Level-3 tasks in Table 2. For Level-1 tasks, both models demonstrated robust fundamental manipulation skills, achieving high success rates due to the simplicity of these single-step operations. In Level-2 tasks, success rates began to decline, particularly for DP. We observed that DP frequently stalled. For instance, in the Heater Beaker task, the DP algorithm halted during the final action until time expired, while in the Operate Drawer task, most DP failures stemmed from its inability to execute any action successfully. In the Stir with GlassRod task, both models exhibited errors in grasping or stirring position due to the small size of the glass rod.
Table 4: Performance comparison of different models when manipulating different size objects.Values represent success rates (%).
Task Level Model Success Rate (%)
ID OOD
Pick ACT 31.2 1.7
DP 11.2 0.0
Pour Liquid ACT 25.6 0.0
DP 8.0 0.0
In Level-3 tasks, the performance gap widened, with ACT maintaining superior stability, while DP struggled with generalization to novel materials and precise actions, such as pressing a button in Heater Beaker. In contrast, ACT showed greater robustness when handling out-of-domain visual features. We tested generalization to out-of-domain shapes in Table 4 by jointly training datasets with objects of varying sizes and evaluating on out-of-domain objects. The results revealed that joint training with objects of significantly different sizes led to a substantial drop in in-domain success rates, with near-zero success rates for out-of-domain shapes and sizes. This suggests that both models largely lack the ability to manipulate out-of-domain objects. Level-4 tasks, as shown in the provided table, assessed long-horizon tasks. ACT significantly outperformed DP, achieving higher success rates in single-stage primitive (SP) actions and sub-steps (A1–A7), though both models experienced sharp declines in later sub-steps (A5–A7) due to cumulative errors in complex sequences. In sub-steps, although each policy is more stable when handling its own task, it is necessary to design specific action decomposition and switching mechanisms, which increases the overall system complexity. Moreover, the transitions between actions may become discontinuous or suffer from distribution shift issues.
ACT exhibited greater stability across most subsequent tasks. We hypothesize that this disparity arises from several factors. First, compared to ACT, DP’s shorter prediction horizon makes it more prone to stagnation. For example, in the Heater Beaker task, DP often hovered above the button without pressing it in the final step. Our task setting outputs joint position, and DP’s outputs are more jittery than ACT’s, increasing the likelihood of dropping grasped objects. Second, DP’s high sensitivity to visual features led to inconsistent performance, particularly in the Stir with GlassRod task, where its output struggled to accurately locate the glass rod. Furthermore, in generalization tests for Level-3 tasks, DP experienced significant performance declines when encountering unseen materials. These findings suggest that future work should focus on enhancing the models’ ability to generalize across diverse object properties and task variations. Please refer to the Supplementary Materials for more details.
6 Conclusion and Limitations
We propose a simulation and evaluation platform for embodied agents in scientific laboratory settings. To this end, we have designed a hierarchical benchmark covering five levels and more than 30 tasks, and constructed a high-fidelity simulation environment for physical and chemical processes, along with a rich library of laboratory assets. We conducted a systematic evaluation of mainstream imitation learning methods, and the results show that current imitation learning approaches still exhibit significant shortcomings when tackling complex long-horizon tasks. While these methods demonstrate some degree of visual generalization, they show little to no generalization to objects with different shapes. Overall, these findings indicate that, despite notable progress in embodied agents and their underlying models, achieving general-purpose agents for scientific laboratory scenarios will require continued in-depth research and innovation. While LabUtopia provides a scalable and versatile simulation platform, several limitations remain: i) A primary limitation is that the current benchmark operates in simulation. Bridging the sim-to-real gap by building a physical environment is a crucial direction for future work. ii) LabUtopia supports two robot embodiments (Fetch and Panda). Expanding support to a broader range of robots would enable the collection of more heterogeneous action trajectories and facilitate research on generalization. iii) In model selection, we failed to test vision-language-action (VLA) models, which may limit the exploration and evaluation of more advanced multimodal learning approaches, particularly in complex laboratory tasks requiring the integration of vision, language, and action.